OpenStudy (anonymous):

The graph of an equation contains the following three points: (11, 64), (22, 72), and (33, 81). Is the graph linear or nonlinear?

OpenStudy (anonymous):

\[\frac{72-64}{22-11}=\frac{81-72}{33-22}\] ? \[\frac{8}{11}\neq \frac{9}{11} \] Nonlinear, the slopes are not equal.